OpenStudy (anonymous):

0.5(1,3) 1.7(2,5) 1.7(1,3) 0.5(1,3)

OpenStudy (amistre64):

i take it those are vectors

OpenStudy (amistre64):

\[proj_{v}u=\frac{u.v}{|v|^2}v\]

OpenStudy (amistre64):

(2+15)/sqrt(10) looks to be the scalar

OpenStudy (amistre64):

lol, make that a 10 :)

OpenStudy (amistre64):

15+2 = 17 17/10 = 1.7